Meck Company is considering the acquisition of another entity. The following data relate to the acquiree:


Shareholders’ equity 5,000,000

Earnings for prior three years 1,500,000


The acquiree has a valuable patent which is not recorded. If the entity is sold, the patent would be transferred to the buyer for P500,000. Other assets are properly appraised. The patent has a remaining life of 5 years. The earnings of the entity are expected to increase 10% more than the average earnings of the past three years before taking into consideration the amortization of the patent cost.


Compute for the goodwill under the following methods:

1. Average future earnings are capitalized at 8%

2. Goodwill is measured at the average excess earnings capitalized at 10% with normal rate at 8%.

3. Goodwill is measured at the present value of the average excess earnings discounted at 10% for four years with normal rate of 8%. The present value of an ordinary annuity of 1 for 4 years at 10% is 3.17.


1
Expert's answer
2021-02-16T05:53:25-0500
"Solution"

"1. FV=PV(1+r)^n\\
500,000=(1+0.1)^5\\
500,000(1.6105)\\
=805250\\
Earnings=\frac{1.500,000}{3}\\
=500,000\ p.a\\
10\%\ increase=\frac{110}{100}\times500,000=550,000p.a\\
5years=5\times550,000\\
=2,750,000\\
Goodwill=Average\ profits\times\ no\ of\ years\\
Average\ profit=\frac{profit}{output}\\
=\frac{2,750,000}{275}=10,000"

"2. Average\ excess\ earnings=500,000\times\frac{110}{100}\\
550,000\\
Excess=550,000-50,000\\
=50,000\\
Goodwill=50,000\times5=250,000\\
=\frac{250,000}{27000}\\
=1000""3.PV=500,000\\
Average\ excess\ earnings\\
=5000000\times5\\
=25000000-250000\\
after discounting=2475000=\frac{8}{(1.1)^4}\\
13,5236.6"